OpenStudy (anonymous):

How do you find the exact solution of: log(x+3)+log(x)=1

OpenStudy (anonymous):

log(x^2+3x)=1 x^2+3x=e x^2+3x-e=0

OpenStudy (anonymous):

did u like the answer?

OpenStudy (anonymous):

I don't really understand it actually. I don't know what the e is..

OpenStudy (anonymous):

e is an exponential number e=2.7 (approx.)

OpenStudy (anonymous):

Ok so the answer is that last line or zero?

OpenStudy (anonymous):

zero ??? use the quadratic formula to solve for x... what grade r u in?